OpenStudy (anonymous):

Simplify 5 to the negative 4th power over 5 to the 3rd. A 5*7 B 5*−1 C 1 over 5 D 1 over 5 to the 7th power

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

To divide powers with the same base, give the same base, and subtract the exponents.

OpenStudy (mathstudent55):

Example: \(\dfrac{8^{-5}}{8^2} = 8^{-5 - 2} = 8^{-7} \)

OpenStudy (skullpatrol):